A complex is represented as CoCl3.xNH3 It`s 0.1 m solution in aqueous solution shows ∆T=0.558C Of(H2O)=1.86/molK and assume 100% ionisation and Co ordination number of Co(|||) is 6.what is the complex?

Grade:12

1 Answers

Vikas TU
7 years ago
From Tf = i*kf*m
0.558 = i*1.86*0.1
i => 0.558/(1.86*0.1) = 3
Therfore,
x = 3 as the CN. is = 6
to fulfill it Total ions are 3Cl- and 3NH3